OpenStudy (anonymous):

what is 1 2/5 times 4 2/3

OpenStudy (texaschic101):

you first need to turn your mixed numbers into improper fractions, then just multiply. 1 2/5 = (1 * 5 + 2)/5 = 7/5 4 2/3 = (4 * 3 + 2)/3 = 14/3 now we have : 7/5 x 14/3 = 98/15 and if you need it back into a mixed number, then it is 6 8/15

OpenStudy (anonymous):

6 8/15 I did it on a calculator

OpenStudy (skyz):

6 8/15 Correct :)